OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(6, 1) with slope 9.

OpenStudy (anonymous):

Use the point slope formula: y - y1 = m(x-x1)

OpenStudy (anonymous):

m = 9, x1 = 6, y1 = 1

OpenStudy (anonymous):

y-y1 = m(x-x1) y-1= 9(x-6) y-1 = 9x-54 y= 9x-53

OpenStudy (anonymous):

y-1 = 9(x-6) This is in point-slope form

OpenStudy (anonymous):

@study100 that would be slope-intercept form. The question is asking for point-slope form

OpenStudy (anonymous):

A yes then that would be y-1=9(x-6) ^^

OpenStudy (anonymous):

Thank you